A particle undergoes simple harmonic motion with amplitude 25 cm and maximum speed 4.8 m/s. If a stopwatch is started when the particle is pulled to its maximum position and released, what is the formula for the position of the particle as a function of time?
Answer #1

angular frequency

w = v/ A = 4.8 / 0.25 = 19.2 rad/s.

now as we know

x = A cos ( wt)

x = 25 cos (19.2 t)

======

Comment before rate in case any doubt, will reply for sure.. goodluck